OpenStudy (anonymous):

The celestial sphere shown at right has radius 9 inches. The planet in the sphere’s center has radius 3 inches. What is the ratio of the volume of the planet to the volume of the celestial sphere? What is the ratio of the surface area of the planet to the surface area of the celestial sphere?

OpenStudy (anonymous):

What is the formula for the volume of a sphere?

OpenStudy (anonymous):

4/3 pi3 right

OpenStudy (anonymous):

4/3 pi r^3

OpenStudy (anonymous):

well find the volume of each one first where the first volume has r=9 and the second has r=3 Why do we want the volume? Well we want to compare them right.

OpenStudy (anonymous):

i have for r=9 is 972pi3

OpenStudy (anonymous):

and r=3 is 36pi3

OpenStudy (anonymous):

Looks right but Im wondering where the 3 came from :P in pi3 I think you put it in by mistake.

OpenStudy (anonymous):

you know what ratio is right? basically a fraction and fractions is when you divide. You have to values so you divide one value with the other and you have the ratio

OpenStudy (anonymous):

so that mean 1:3 right?

OpenStudy (anonymous):

is \[\frac{972\pi}{36\pi}\] 1:3?

OpenStudy (anonymous):

so i got 27

OpenStudy (anonymous):

if tha's what is then yes for the volume. Do the same for surface area. Find the formula for surface area and use r=9 and r=3

OpenStudy (anonymous):

324pi and 36pi